A 200m long is cut into two parts such that one part is 12cm longer than the other.how long is each part?
Answer by checkley79(3341)

LET X=THE SHORTEST LENGTH
LET (200-X) BE THE LONGEST LENGTH.
(200-X)=X+12
-X-X=12-200
-2X=-188
X=-188/-2
X=94 INCHES FOR THE SHORTEST LENGTH.
200-94=106 INCHES FOR THE LONGEST LENGTH.
PROOF:
94+106=200
200=200